# Maker-Breaker domination game played on corona products of graphs

### v1

@@ Abstract
${}^{a}$ Department of Mathematics, Mar Athanasius College, Kothamangalam, India
${}^{b}$ Department of Mathematics, Pavanatma College, Murickassery, India
${}^{c}$ Faculty of Mathematics and Physics, University of Ljubljana, Slovenia
${}^{d}$ Institute of Mathematics, Physics and Mechanics, Ljubljana, Slovenia
${}^{e}$ Faculty of Natural Sciences and Mathematics, University of Maribor, Slovenia
@@ 1 Introduction
The following graph invariants are naturally associated to the MBD game [4, 15]. The Maker-Breaker domination number, $\gamma_{\rm MB}(G)$, is the minimum number of moves of Dominator to win the D-game on $G$ when both players play optimally. That is, $\gamma_{\rm MB}(G)$ is the minimum number of moves of Dominator such that he wins in this number of moves no matter how Staler is playing. If Dominator has no winning strategy in the D-game, then set $\gamma_{\rm MB}(G)=\infty$. The Staller-Maker-Breaker domination number, $\gamma_{\rm SMB}(G)$, is the minimum number of moves of Staller to win the D-game on $G$ when both players play optimally, where $\gamma_{\rm SMB}(G)=\infty$ if Staller has no winning strategy. In a similar manner, $\gamma_{\rm MB}^{\prime}(G)$ and $\gamma_{\rm SMB}^{\prime}(G)$ are the two parameters associated with the $S$-game. Briefly, we will call $\gamma_{\rm MB}(G)$, $\gamma_{\rm MB}^{\prime}(G)$, $\gamma_{\rm SMB}(G)$, and $\gamma_{\rm SMB}^{\prime}(G)$ the MBD numbers of $G$.
In the seminal paper [9] it was proved, among other things, that deciding the winner of the MBD game can be solved efficiently on trees, but it is PSPACE-complete in general. These results have been very recently strengthened in [10] by proving that the game is PSPACE-complete on split and bipartite graphs and by developing a linear time algorithm to solve this game in forests. The paper [4] focuses on $\gamma_{\rm SMB}(G)$ and $\gamma_{\rm SMB}^{\prime}(G)$ and among other results establishes appealing exact formulas for $\gamma_{\rm SMB}^{\prime}(G)$ where $G$ is a path. In [5], for every positive integer $k$, trees $T$ with $\gamma_{\rm SMB}(T)=k$ are characterized and exact formulas for $\gamma_{\rm SMB}(G)$ and $\gamma_{\rm SMB}^{\prime}(G)$ derived for caterpillars. In the main result of [13], $\gamma_{\rm MB}$ and $\gamma_{\rm MB}^{\prime}$ are determined for Cartesian products of $K_{2}$ by a path. In [7], the MBD game is further studied on Cartesian products of paths, stars, and complete bipartite graphs. The total version of the MBD game was introduced in [14] and further investigated in [12].
@@
Let $G=(V(G),E(G))$ be a graph. The order of $G$ will be denoted by $n(G)$. For a vertex $v\in V(G)$, its open neighbourhood is denoted by $N(v)$ and its closed neighbourhood by $N[v]$. The degree of $v$ is $\deg(v)=|N(v)|$. The minimum and the maximum degree of $G$ are denoted by $\delta(G)$ and $\Delta(G)$. An isolated vertex is a vertex of degree 00, a leaf is a vertex of degree $1$. A support vertex is a vertex adjacent to a leaf, a strong support vertex is a vertex that is adjacent to at least two leaves. A dominating set of $G$ is a set $D\subseteq V(G)$ such that each vertex from $V(G)\setminus D$ has a neighbour in $D$. The domination number $\gamma(G)$ of $G$ is the minimum cardinality of a dominating set of $G$. A vertex of degree $n(G)-1$ is a dominating vertex.
@@ 2 Outcome of the game on corona products
Therefore, we may indeed restrict in the rest of the paper to corona products $G\odot H$, where $G$ is a connected graph with $n(G)\geq 2$. In what follows, we will also always assume that $V(G)=\{v_{1},\ldots,v_{n(G)}\}\subseteq V(G\odot H)$, and denote by $H_{i}$ the copy of $H$ in $G\odot H$ assigned to $v_{i}$, $i\in[n(G)]$.
@@
If $G$ is a connected graph with $n(G)\geq 2$ and $H$ is a graph, then
@@ 3 MBD numbers of corona products
If $o(H)=\mathcal{S}$ and $G$ is a connected graph with $n(G)\geq 2$, then
@@
If $H\in\mathcal{D}$ and $G$ is a connected graph with $n(G)\geq 2$, then
@@
Following the above strategy, Dominator will select at least $\left\lceil\frac{n(G)}{2}\right\rceil$ vertices of $G$. If Staller was the first to select a vertex in $H_{i}$, then Dominator will select at most at most $\gamma_{\rm MB}^{\prime}(H)$ vertices of $H_{i}$. Similarly, If Dominator was the first to select a vertex in $H_{i}$, then he will select at most at most $\gamma_{\rm MB}(H)$ vertices of $H_{i}$. Since $\gamma_{\rm MB}(H)\leq\gamma_{\rm MB}^{\prime}(H)$ (see [15, Theorem 3.1]), Dominator will play at most $\gamma_{\rm MB}^{\prime}(H)$ vertices of each $H_{i}$, where $v_{i}$ was selected by Staller. As there are at most $\left\lfloor\frac{n(G)}{2}\right\rfloor$ such vertices, the upper bound for $\gamma_{\rm MB}(G\odot H)$ follows.
@@
If $H\in\mathcal{D}$ with $\gamma_{\rm MB}(H)=\gamma_{\rm MB}^{\prime}(H)$, and $G$ is a connected graph with $n(G)\geq 2$, then
@@
If $H\in\mathcal{D}$, then $\gamma_{\rm MB}(K_{2}\odot H)=1+\gamma_{\rm MB}(H)$.
@@ 4 Corona products with small MBD numbers
[14, Proposition 3.2] Let $G$ be a graph with $\gamma(G)=2$. Then $\gamma_{\rm MB}(G)=\gamma(G)=2$ if and only if $G$ has a vertex that lies in at least two $\gamma$-sets of $G$.
We next complement Proposition 4.1 with the following result.
@@
Let $G$ be a graph with $\gamma(G)\leq 2$. Then $\gamma_{\rm MB}^{\prime}(G)=2$ if and only if $G$ has at most one dominating vertex and at least two vertices that lie in two different dominating sets of $G$ of cardinality $2$.
Proof. Assume first that $\gamma_{\rm MB}^{\prime}(G)=2$. Then $G$ has at most one dominating vertex for otherwise Dominator can finish the game in his first move. Suppose that $G$ contains at most one vertex that lies in two different dominating sets of $G$. If $x$ is a such a vertex, Staller can select $x$ as her first move. Then, no matter how Dominator plays, he needs at least three moves to finish the game.
Conversely, assume that $G$ has at most one dominating vertex and at least two vertices that lie in two different dominating sets of $G$ of cardinality $2$. Then Staller can prevent Dominator from winning in one move by selecting first a possible dominating vertex. However, no matter how Staller starts, after her move there exists a not yet played vertex which lies in two dominating sets of $G$ of cardinality $2$. By selecting such a vertex, Dominator will win the game in his second move. $\square$
@@
To describe graphs with SMB numbers equal to three is more tricky. Nevertheless we can do it for $\gamma_{\rm SMB}^{\prime}$ as follows.
@@
    There exists a vertex $u\in V(G)$ such that either $G-u$ has at least two strong support vertices or $G-u$ has exactly one isolated vertex and a strong support vertex.
Proof. Assume first that $\gamma_{\rm SMB}^{\prime}(G)=3$. Then $G$ has no isolated vertices and no strong support vertices. Let $u$ be an optimal first move of Staller in the S-game. After this move, the remaining game is a D-game on $G-u$ and we have $\gamma_{\rm SMB}(G-u)=2$. In view of Proposition 4.4 this is possible only when either $G-u$ has at least two strong support vertices or $G-u$ has exactly one isolated vertex and a strong support vertex.
Conversely, assume that $G$ is a graph which satisfies (i) and (ii). Let $u$ be a vertex that satisfies (ii) and let $u$ be the first move of Staller in the S-game. If $G-u$ has an isolated vertex then Dominator selects this vertex for otherwise Staller could claim it in the next move to win the game. After that, Staller selects a strong support vertex and wins the game in her third move. Similarly, if $G-u$ has no isolated vertex, then $G-u$ has two strong support vertices and again Staller can win in her third move. We conclude that $\gamma_{\rm SMB}^{\prime}(G)=3$. $\square$
For corona products Proposition 4.6 yields:
Corollary 4.7
An alternative way to derive Corollary 4.7 is to use Theorem 3.1, that is, $\gamma_{\rm SMB}(G\odot H)=\gamma_{\rm SMB}^{\prime}(G\odot H)=3$ if and only if $\gamma_{\rm SMB}(H)=2$. If $\delta(H)\geq 1$, then Proposition 4.4(i) implies that this is if and only if $H$ has at least two strong support vertices. On the other hand, if $H$ has an isolated vertex, we infer that it must at least one strong support vertex.
@@
Let $G$ be a graph with $n(G)\geq 2$ and let $H$ be graphs. Then $\gamma_{\rm MB}(G\odot H)=3$ if and only if either $n(G)=2$ and $\gamma_{\rm MB}(H)=2$, or $n(G)=3$ and $H$ has a dominating vertex.
@@ 5 Concluding remarks
In this paper, we explore the Maker-Breaker domination game on corona products of graphs. We have identified some general features, but several directions remain open for further exploration. Let us point to some of them. In Theorem 3.4 we have determined $\gamma_{\rm MB}(G\odot T)$ and $\gamma_{\rm MB}^{\prime}(G\odot T)$ for an arbitrary graph $G$ with with $n(G)\geq 2$ and an arbitrary tree $T$ with a perfect matching. It would be interesting to derive such a result also for arbitrary trees, that is, also for trees which do not contain perfect matchings. With respect to small MBD numbers it in particular remains to consider arbitrary graphs $G$ with $\gamma_{\rm SMB}(G)=3$, or with $\gamma_{\rm MB}(G)=3$, or with $\gamma_{\rm MB}^{\prime}(G)=3$, as well as corona products $G\odot H$ with $\gamma_{\rm MB}^{\prime}(G\odot H)=3$.
@@ Acknowledgements
Sandi Klavžar was supported by the Slovenian Research Agency ARIS (research core funding P1-0297 and projects J1-2452, N1-0285).

### v2

@@ Abstract
a Department of Mathematics, Mar Athanasius College, Kothamangalam, India
b Department of Mathematics, Pavanatma College, Murickassery, India
c Faculty of Mathematics and Physics, University of Ljubljana, Slovenia
d Institute of Mathematics, Physics and Mechanics, Ljubljana, Slovenia
e Faculty of Natural Sciences and Mathematics, University of Maribor, Slovenia
@@ 1 Introduction
The following graph invariants are naturally associated to the MBD game [4, 16]. The Maker-Breaker domination number, $\gamma_{\rm MB}(G)$, is the minimum number of moves of Dominator to win the D-game on $G$ when both players play optimally. That is, $\gamma_{\rm MB}(G)$ is the minimum number of moves of Dominator such that he wins in this number of moves no matter how Staller is playing. If Dominator has no winning strategy in the D-game, then set $\gamma_{\rm MB}(G)=\infty$. The Staller-Maker-Breaker domination number, $\gamma_{\rm SMB}(G)$, is the minimum number of moves of Staller to win the D-game on $G$ when both players play optimally, where $\gamma_{\rm SMB}(G)=\infty$ if Staller has no winning strategy. In a similar manner, $\gamma_{\rm MB}^{\prime}(G)$ and $\gamma_{\rm SMB}^{\prime}(G)$ are the two parameters associated with the S-game. Briefly, we will call $\gamma_{\rm MB}(G)$, $\gamma_{\rm MB}^{\prime}(G)$, $\gamma_{\rm SMB}(G)$, and $\gamma_{\rm SMB}^{\prime}(G)$ the MBD numbers of $G$.
In the seminal paper [10] it was proved, among other things, that deciding the winner of the MBD game can be solved efficiently on trees, but it is PSPACE-complete even for bipartite graphs and split graphs. In [11] the authors give a complex linear algorithm for the Maker-Maker version of the game. The paper [4] focuses on $\gamma_{\rm SMB}(G)$ and $\gamma_{\rm SMB}^{\prime}(G)$ and among other results establishes appealing exact formulas for $\gamma_{\rm SMB}^{\prime}(G)$ where $G$ is a path. In [5], for every positive integer $k$, trees $T$ with $\gamma_{\rm SMB}(T)=k$ are characterized and exact formulas for $\gamma_{\rm SMB}(G)$ and $\gamma_{\rm SMB}^{\prime}(G)$ derived for caterpillars. In the main result of [14], $\gamma_{\rm MB}$ and $\gamma_{\rm MB}^{\prime}$ are determined for Cartesian products of $K_{2}$ by a path. In [8], the MBD game is further studied on Cartesian products of paths, stars, and complete bipartite graphs. The total version of the MBD game was introduced in [15] and further investigated in [13].
@@
Let $G=(V(G),E(G))$ be a graph. The order of $G$ will be denoted by $n(G)$. For a vertex $v\in V(G)$, its open neighbourhood is denoted by $N(v)$ and its closed neighbourhood by $N[v]$. The degree of $v$ is $\deg(v)=|N(v)|$. The minimum and the maximum degree of $G$ are denoted by $\delta(G)$ and $\Delta(G)$. An isolated vertex is a vertex of degree 00, a leaf is a vertex of degree $1$. A support vertex is a vertex adjacent to a leaf, a strong support vertex is a vertex that is adjacent to at least two leaves. A dominating set of $G$ is a set $D\subseteq V(G)$ such that each vertex from $V(G)\setminus D$ has a neighbour in $D$. The domination number $\gamma(G)$ of $G$ is the minimum cardinality of a dominating set of $G$. A dominating set of $G$ of cardinality $\gamma(G)$ is refered to as a $\gamma$-set of $G$. A vertex of degree $n(G)-1$ is a dominating vertex.
@@
For $X\subseteq V(G)$, let $G|X$ denote the graph $G$ in which vertices from $X$ are considered as being already dominated.
@@ 2 Outcome of the game on corona products
Therefore, we may indeed restrict in the rest of the paper to corona products $G\odot H$, where $G$ is a graph with $n(G)\geq 2$. In what follows, we will also always assume that $V(G)=\{v_{1},\ldots,v_{n(G)}\}\subseteq V(G\odot H)$, and denote by $H_{i}$ the copy of $H$ in $G\odot H$ assigned to $v_{i}$, $i\in[n(G)]$.
@@
If $G$ is a graph with $n(G)\geq 2$ and $H$ is a graph, then
@@ 3 MBD numbers of corona products
If $o(H)=\mathcal{S}$ and $G$ is a graph with $n(G)\geq 2$, then
@@
If $o(H)=\mathcal{D}$ and $G$ is a graph with $n(G)\geq 2$, then
@@
Following the above strategy, Dominator will select at least $\left\lceil\frac{n(G)}{2}\right\rceil$ vertices of $G$. If Staller was the first to select a vertex in $H_{i}$, then Dominator will select at most at most $\gamma_{\rm MB}^{\prime}(H)$ vertices of $H_{i}$. Similarly, If Dominator was the first to select a vertex in $H_{i}$, then he will select at most at most $\gamma_{\rm MB}(H)$ vertices of $H_{i}$. Since $\gamma_{\rm MB}(H)\leq\gamma_{\rm MB}^{\prime}(H)$ (see [16, Theorem 3.1]), Dominator will play at most $\gamma_{\rm MB}^{\prime}(H)$ vertices of each $H_{i}$, where $v_{i}$ was selected by Staller. As there are at most $\left\lfloor\frac{n(G)}{2}\right\rfloor$ such vertices, the upper bound for $\gamma_{\rm MB}(G\odot H)$ follows.
@@
If $o(H)=\mathcal{D}$ with $\gamma_{\rm MB}(H)=\gamma_{\rm MB}^{\prime}(H)$, and $G$ is a graph with $n(G)\geq 2$, then
@@
If $o(H)=\mathcal{D}$, then $\gamma_{\rm MB}(K_{2}\odot H)=1+\gamma_{\rm MB}(H)$.
@@ 4 Corona products with small MBD numbers
[15, Proposition 3.2] Let $G$ be a graph with $\gamma(G)=2$. Then $\gamma_{\rm MB}(G)=2$ if and only if $G$ has a vertex that lies in two $\gamma$-sets of $G$.
If $x\in V(G)$ and $S\subseteq V(G)$, then we say that $S$ is $x$-free if $x\notin S$. We next complement Proposition 4.1 with the following result.
@@
Let $G$ be a graph with $\gamma(G)=2$. Then $\gamma_{\rm MB}^{\prime}(G)=2$ if and only if for any vertex $x\in V(G)$ there exists a vertex $y_{x}\in V(G)$ such that $y_{x}$ lies in two $x$-free $\gamma$-sets of $G$.
Proof. Assume first that $\gamma_{\rm MB}^{\prime}(G)=2$. Let $x$ be an arbitrary vertex of $G$. Then $x$ is a legal first move of Staller. Since $\gamma_{\rm MB}^{\prime}(G)=2$, Dominator has an optimal reply $y$ such that he needs only one more move to finish the game. If $y^{\prime}$ is such a vertex, then $\{y,y^{\prime}\}$ is a $\gamma$-set of $G$. Now Staller can reply by selecting $y^{\prime}$ as her second move. As $\gamma_{\rm MB}^{\prime}(G)=2$, Dominator can finish the game by selecting another vertex $y^{\prime\prime}$. Then $\{y,y^{\prime}\}$ and $\{y,y^{\prime\prime}\}$ are two $\gamma$-sets of $G$ which are clearly $x$-free, hence setting $y_{X}=y$ we have the vertex required.
Conversely, assume that for any vertex $x\in V(G)$ there exists a vertex $y_{x}\in V(G)$ such that $y_{x}$ lies in two $x$-free $\gamma$-sets of $G$. Assume that Staller starts the game by playing an arbitrary vertex $x$ of $G$. Then Dominator replies by selecting $y_{x}$ as her first move. (As $\gamma(G)=2$, the vertex $y_{x}$ is clearly not a dominating vertex.) Since $y_{x}$ lies in two $\gamma$-sets of $G$ (of cardinality $2$), Dominator can win the game in his second move. $\square$
@@
We next describe corona products with both SMB numbers equal to three.
Proposition 4.6
Proof. Recall that Theorem 3.1 asserts $\gamma_{\rm SMB}(G\odot H)=\gamma_{\rm SMB}^{\prime}(G\odot H)=3$ if and only if $\gamma_{\rm SMB}(H)=2$. If $\delta(H)\geq 1$, then Proposition 4.4(i) implies that this is if and only if $H$ has at least two strong support vertices. On the other hand, if $H$ has an isolated vertex, we infer that it must at least one strong support vertex. $\square$
In general, we can characterize the graphs $G$ with $\gamma_{\rm SMB}^{\prime}(G)=3$ as follows.
@@
    There exist a vertex $u\in V(G)$ such that $G-u$ has at least two strong support vertices, or $G-u$ has an isolated vertex and a strong support vertex.
Proof. Let $G$ be a graph with $\delta(G)\geq 1$.
Assume first that $\gamma_{\rm SMB}^{\prime}(G)=3$. Then $G$ has no strong support vertices, for otherwise we would have $\gamma_{\rm SMB}^{\prime}(G)\leq 2$. Let $u$ be an optimal first move of Staller in the S-game.
Claim: After Staller selects $u$, the remaining game is equivalent to the D-game played on $G-u$.
To prove the claim we need to show that $u$ is not the vertex which is undominated (by Dominator) at the end of the game. Assume first that $\deg(u)=1$. If $v$ is the neighbor of $u$, then Dominator must select $v$ in his first move, for otherwise Staller would win the game in her second move which is not possible by our assumption $\gamma_{\rm SMB}^{\prime}(G)=3$. Hence the claim. Assume second that $\deg(u)\geq 2$ and let $v$ be a vertex with smallest degree among the vertices from $N(u)$. Let Dominator select $v$ as his first move. Since $G$ has no strong support vertex, each vertex from $N(u)\setminus\{v\}$ has degree at least $2$, hence Staller cannot finish the game in her second move. This implies that the move $v$ is an optimal move of Dominator which in turn proves the claim.
We have thus seen that after Staller selects $u$, the remaining game is a D-game on $G-u$. As $\gamma_{\rm SMB}^{\prime}(G)=3$, we have $\gamma_{\rm SMB}(G-u)=2$. In view of Proposition  4.4 this is possible only when either $G-u$ has at least two strong support vertices or $G-u$ has exactly one isolated vertex and a strong support vertex.
Conversely, assume that $G$ is a graph which satisfies $(i)$ and $(ii)$. Let $u$ be a vertex that satisfies $(ii)$ and let $u$ be the first move of Staller in the S-game. If $G-u$ has an isolated vertex then Dominator selects this vertex for otherwise Staller could claim it in the next move to win the game. After that, Staller selects a strong support vertex and wins the game in her third move. Similarly, if $G-u$ has no isolated vertex, then $G-u$ has two strong support vertices and again Staller can win in her third move. Thus we conclude that $\gamma_{\rm SMB}^{\prime}(G)=3$. $\square$
@@
Let $G$ be a graph with $n(G)\geq 2$ and let $H$ be a graph. Then $\gamma_{\rm MB}(G\odot H)=3$ if and only if either $n(G)=2$ and $\gamma_{\rm MB}(H)=2$, or $n(G)=3$ and $H$ has a dominating vertex.
@@ 5 Concluding remarks
In this paper, we have explored the Maker-Breaker domination game on corona products $G\odot H$ of graphs. It is worth noting that this game is equivalent to the Maker-Breaker domination game played on the disjoint union of $n(G)$ copies of the graph obtained from $H$ by adding a new vertex adjacent to all the vertices of $H$. In other words, we are interested in the Maker-Breaker domination game played on the disjoint union of several copies of a given graph $H$ with $\gamma(H)=1$. Having this in mind it would be of interest to extend the results of this paper to generalized coronas which are defined just as coronas, except that the graphs associated to vertices of $G$ can be different. Generalized coronas were by now investigated from different angles, cf. [6, 19, 23].
For the the Maker-Breaker domination game on corona products, we have identified some general features, but several directions remain open for further exploration. Let us point to some of them. In Theorem 3.4 we have determined $\gamma_{\rm MB}(G\odot T)$ and $\gamma_{\rm MB}^{\prime}(G\odot T)$ for an arbitrary graph $G$ with with $n(G)\geq 2$ and an arbitrary tree $T$ with a perfect matching. It would be interesting to derive such a result also for arbitrary trees, that is, also for trees which do not contain perfect matchings. With respect to small MBD numbers it in particular remains to consider arbitrary graphs $G$ with $\gamma_{\rm SMB}(G)=3$, or with $\gamma_{\rm MB}(G)=3$, or with $\gamma_{\rm MB}^{\prime}(G)=3$, as well as corona products $G\odot H$ with $\gamma_{\rm MB}^{\prime}(G\odot H)=3$.
@@ Acknowledgements
Sandi Klavžar was supported by the Slovenian Research Agency ARIS (research core funding P1-0297 and projects N1-0218, N1-0285).